OpenStudy (anonymous):

Divide 5x^4 - 2x^2 - 3 by x - 1 Long Division.

OpenStudy (anonymous):

\[5 x^4-2 x^2-3=(x-1) (x+1) \left(5 x^2+3\right) \]

OpenStudy (anonymous):

\[5x^4-2x^2-3 = (5x^2+3)(x+1)(x-1)\]